OpenStudy (anonymous):

Evaluate the function for x = –3. f(x) = 4x – 5 A. 7 B. –17 C. –7 D. 17

OpenStudy (anonymous):

4(-3) - 5 =-12-5 =-17

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Replace x with -3

OpenStudy (mertsj):